OpenStudy (anonymous):

find the quotient and simplify q/7 divided by q/26 help and explain!

OpenStudy (anonymous):

hi diana, do you still need help with this problem or do you think you have it under control?

OpenStudy (anonymous):

I still need help :/ I'm trying to figure it out because any way I do it doesn't seem to fit one of the answers

OpenStudy (anonymous):

Its multiple choice

OpenStudy (anonymous):

I think its \[\frac{ q ^{2} }{ 182}\]

OpenStudy (anonymous):

:) no worries. I have a question. How would you do the problem (1/2)/(1/4)?

OpenStudy (anonymous):

times it across and you get 2/4 = 1/2?

OpenStudy (anonymous):

division of fractions means that you take the fraction you are dividing by, invert it (flip it), and then multiply straight across.

OpenStudy (anonymous):

so, in the case of 1/2 divided by 1/4, you would flip the dividing fraction (1/4) to 4/1 and you would get \[\frac{ 1 }{ 2 }\times \frac{ 4 }{ 1 }\]

OpenStudy (anonymous):

then multiply straight across to get 4/2

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so answer is 2?

OpenStudy (anonymous):

okeedokee, this then simplifies to 2

OpenStudy (anonymous):

perfect. yes.

OpenStudy (anonymous):

now in the case of the problem you are doing, what is the dividing fraction?

OpenStudy (anonymous):

so \[\frac{ q }{ 7 } \div \frac{ 27 }{ q } then you multiply?\]

OpenStudy (anonymous):

then you multiply*

OpenStudy (anonymous):

Imean 26******

OpenStudy (anonymous):

close, you invert the fraction. that changes it from a division problem to a multiplication problem

OpenStudy (anonymous):

so \[\frac{ q }{ 182 }\]

OpenStudy (anonymous):

\[\frac{ q }{ 7 }\times \frac{ 26 }{ q }\]

OpenStudy (anonymous):

do you see how i flipped the dividing fraction and then it is no longer division but multiplication?

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

so what is your final answer?

OpenStudy (anonymous):

notice that the q's cross cancel

OpenStudy (anonymous):

\[\frac{ q ^{2} }{ 182 }\] ?

OpenStudy (anonymous):

thats the closet thing to my multple choice

OpenStudy (anonymous):

are you familiar with cross cancellation in the product of two fractions? so, for instance, if you are multiplying \[\frac{ 2 }{ 3 }by \frac{ 1 }{ 2 }\], you will be able to cancel the 2 twos and multiply straight across. this would give an answer of 1/3

OpenStudy (anonymous):

now think about your problem and the product above

OpenStudy (anonymous):

yeah , do you want me to write out the multiple choice?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

just the answers.

OpenStudy (anonymous):

\[A. \frac{ 7 }{ 26 }\]

OpenStudy (anonymous):

B.\[B. \frac{ 26 }{ 7 }\]

OpenStudy (anonymous):

\[C. \frac{ q ^{2} }{ 182 }\]

OpenStudy (anonymous):

\[D.\frac{ 26q ^{2} }{ 7 }\]

OpenStudy (anonymous):

ok diana, the key concept here is in the cancellation of the q's.

OpenStudy (anonymous):

so its 7/26

OpenStudy (anonymous):

omg this is hard -_-

OpenStudy (anonymous):

remember, that when i inverted the dividing fraction that the problem became \[\frac{ q }{ 7 }\times \frac{ 26 }{ q }\]

OpenStudy (anonymous):

do remember how to cross cancel?

OpenStudy (anonymous):

it is not hard when you see how to do it... you will be fine.

OpenStudy (anonymous):

yes cancel out the q so it 26/7 ?

OpenStudy (anonymous):

yess!!!!!

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

YAY THANKS!!

OpenStudy (anonymous):

so the steps of dividing by a fraction are: 1. flip the dividing fraction

OpenStudy (anonymous):

2. cross cancel if possible

OpenStudy (anonymous):

2. change the problem after the inversion to multiplication. sorry, do this BEFORE cross cancelling

OpenStudy (anonymous):

then cross cancel. then multiply straight across. it is formulaic when you get used to it

OpenStudy (anonymous):

and remember, it works for ALL fractions, even those with variables. ok?

OpenStudy (anonymous):

i think you are through the worst of it. :). nice job of hangin in.

OpenStudy (anonymous):

any questions? anything still hazy or confusing?

OpenStudy (anonymous):

nope thanks alot

OpenStudy (anonymous):

my pleasure. you take care diana.

OpenStudy (anonymous):

you too (:

OpenStudy (anonymous):

:) bye